# ResearchTown: Simulator of human research community

### v1

@@ 9 Ablation study: ResearchTown is robust
We conduct ablation studies on the number of neighborhood nodes during aggregation. Results indicate that ResearchTown can robustly perform high-quality simulation under both given various numbers of neighborhood nodes. More details are available in Appendix §E.
@@
Ablation on paper writing. When simulating non-existent papers with ResearchTown, users can assign random papers and researchers, making robustness to the number of papers and researchers critical. As shown in Figure 4, papers from the related work section have the greatest impact, increasing the similarity score from 67.3 to 67.5 compared to using all papers. In contrast, using cited papers from the introduction section reduces the score to 65.9, while papers from other sections lower it further to 58.9. Similarly, as illustrated in Figure 6, under agent-aggregation settings, increasing the number of researchers consistently enhances simulation quality. The most notable improvement occurs when increasing from 1 to 2 researchers, boosting the similarity score from 47.6 to 51.5.
Ablation on review writing. We analyze the impact of reviewer numbers on review writing, as reviewers are the only freely selectable factor when using ResearchTown to simulate review writing, making robustness to this factor crucial. As shown in Figure 6, increasing the number of reviewers consistently enhances review quality. Strength scores improve from 65.1 to 66.0, and weakness scores increase from 61.0 to 61.4 when moving from 1 to 5 reviewers.
10 Case study: Out-of-distribution ResearchTown application
As described in Section §8, the node masking-based evaluation of ResearchTown focuses on in-distribution cases, where predefined neighborhoods are available for each data point. However, in real-world use, ResearchTown is designed to generate non-existent papers and reviews without pre-defined neighborhoods, requiring the automatic construction of neighborhoods by matching researchers and papers. This process introduces out-of-distribution scenarios, such as interdisciplinary research, where unrelated papers and researchers from distinct domains form an unconventional neighborhood for a target paper node, with no similar related works existing beforehand.
ResearchTown can inspire interdisciplinary research. Interdisciplinary research is often difficult to conduct in the real world due to a lack of collaboration between researchers from different fields. ResearchTown overcomes this barrier by enabling researcher agents with diverse expertise to read papers, collaborate, and generate innovative ideas together. For instance, as shown in Figure 7, combining researchers and papers from NLP and astronomy produces the novel idea of using kinematic models to study the evolution of linguistic systems. Similarly, combining researchers and papers from NLP and criminology led to the concept of leveraging LLMs to support communities impacted by mass incarceration. We find that there is little existing work combining such domains. These concrete examples highlight how ResearchTown facilitates the creation of pioneering and unexplored cross-domain research ideas.
ResearchTown-written contents might have limited use in the real world. We observe failure modes in ResearchTown when combining researchers and papers from too many domains. The generated content often becomes a meaningless mix of terms. For example, combining researchers and papers from LLM, biology, criminology, and astronomy, ResearchTown generates a research question of “How does coded language in political discourse influence societal biases, and how can a Bayesian hierarchical model be employed to analyze this effect while simultaneously addressing observational biases in white dwarf population studies?” It simply strings together terminology from different domains without presenting a clear research direction. Such vagueness in implementation and analysis details might hinder the real use of the papers simulated from ResearchTown.
@@ 10 Case study: Out-of-distribution ResearchTown application
Figure 7: Examples of generated interdisciplinary research papers from ResearchTown. For each example, we include ResearchTown’s responses to two questions: “What is the problem?” and “What are the key components of my approach and results?” as these are the most critical among the five questions mentioned in Section §6. Appendix §H provides the full contents of the above two and more examples for interdisciplinary research.
@@ 11 Conclusion
We propose ResearchTown as a graph-based multi-agent framework for research community simulation. We start by defining an agent-data graph as an abstract model to describe the human research community. Furthermore, we define a TextGNN framework that describes the message-passing process on the agent-data graph. By treating the research community as a specialized form of the agent-data graph, ResearchTown unifies various research activities—including paper reading, paper writing, and review writing—into a cohesive inference process powered by TextGNN. With the help of ResearchTown, we can simulate research activities that closely mirror real-world efforts and leverage multiple agents to enhance the robustness of the simulation. Additionally, ResearchTown further inspires interdisciplinary research, which is rare in the real world, allowing research agents from different domains to brainstorm novel ideas together. We believe that ResearchTown provides a useful platform and opens new avenues for systematic evaluation and the development of algorithms for automating research processes.
@@ Appendix A Individual contribution
Kunlun Zhu Co-lead, benchmark collection, code writing, paper writing experiment
@@ Appendix B Ethical concern
The development and deployment of ResearchTown raises several important ethical considerations that we have carefully addressed in our work. We first discuss how ResearchTown prevents dangerous use including facilitating plagiarism, producing misleading or low-quality claims, and role-playing human researchers. Furthermore, we discuss the attribution and authorship issues for generated content and discuss the model and data license in our work.
@@ B.1 Potential to facilitate plagiarism
Generative AI’s capabilities for image and text generation can potentially lead to plagiarism in research (Elali & Rachid, 2023). To address this, we have implemented safeguards to ensure responsible usage. ResearchTown is designed as an assistive tool to help researchers gather inspiration for paper and review writing, rather than generating complete, ready-to-use content. By design, ResearchTown ensures that its outputs serve as a starting point for further intellectual effort, rather than a replacement for human researchers.
@@ B.2 Potential to produce misleading or low-quality claims
The motivation behind our paper is to simulate research activities and generate preliminary research progress (e.g., papers and reviews) that can be scrutinized and validated by human researchers, ultimately contributing to the acceleration of the research process. We acknowledge that AI-generated ideas may vary in quality, and therefore, these outputs are not intended for direct dissemination. Instead, they serve as initial, unofficial suggestions that require further experimentally validated by human researchers. This approach ensures that only rigorously tested and verified research is presented as final, high-quality work.
@@ B.4 Attribution and authorship
B.5 Data license
All papers in ResearchBench come from top-tier machine learning conferences (ICLR 2024 and NeurIPS 2024). These papers are publicly available and under the license of CC-BY 4.0, allowing for redistribution and sharing. For the evaluation results of ResearchBench, all inputs and outputs are logged and open for access. Additionally, we keep an accessible record of all supplementary papers referenced during ResearchTown’s inference process. All outputs from ResearchTown are released by the licenses of the papers used for generation.
B.6 Model license
Our work relies solely on the GPT-4o mini, text-embedding-large-3, and voyage-3 models for practical use. Specifically, we reference the model version gpt-4o-mini-2024-07-18 accessed via its API. All of these three models are closed-sourced and operate under a proprietary license. Since we run on public data only and our work is for non-commercial and academic use, we do not violate any restrictions.
Appendix C ResearchTown technical details
In this section, we provide more explanation and implementation details towards ResearchTown simulation algorithm. To achieve better performance and efficiency, we design different prompts for each agent function and make the aggregation process run in parallel.
C.1 ResearchTown aggregation setting implementation
@@
Self-agg. We do not rely on any neighborhood information.
Paper writing: The LLM agent without profiles brainstorms independently without referencing any external data or other agents’ ideas. This setting extends Equation 6 to
@@
Review writing: The LLM agent without profiles writes the review based solely on the paper itself, without considering additional references or other agents. This setting extends Equation 7 to
@@
Agent-agg. We rely only on agent nodes and exclude data nodes.
Paper writing: Multiple LLM agents collaborate by sharing their content and insights to produce the final paper’s content. This setting extends Equation 6 to
@@
Review writing: Multiple LLM agents collectively review the paper, sharing their input and critiques to form the final review. This setting extends Equation 7 to
@@
Data-agg. We rely only on data nodes and exclude agent nodes.
Paper writing: A single LLM agent without profiles reads and synthesizes information from related data sources to write a paper. This setting extends Equation 6 to
@@
Review writing: A single LLM agent without profiles produces a review by reading both the paper and its related data sources, integrating the information to form a comprehensive critique. This setting extends Equation 7 to
@@
Global-agg. We include all neighborhood nodes (both agent and data) during aggregation.
Paper writing: Multiple LLM agents produce content in parallel while referencing various data sources. Their aggregated outputs, which incorporate insights from both other agents and data, form the final paper. This setting extends Equation 6 to
@@
Review writing: Multiple LLM agents each consider the paper and its related works to write their review. The final review is a combination of these integrated perspectives. This setting extends Equation 7 to
@@
C.2 ResearchTown simulation algorithm implementation
@@
C.3 Agent function implementation
@@
Appendix D ResearchBench technical details
In this section, we provide the technical details included in the construction process of ResearchBench. We describe the methodologies used for data collection across its three main components and we name them as: (1) PaperBench, (2) HighImpactPaperBench, and (3) ReviewBench. Statistically, PaperBench and HighImpactPaperBench focus on a paper writing simulation, which contains 1,000 and 100 tasks, respectively. ReviewBench focuses on review writing simulation and includes 200 tasks.
D.1 Data collection details

### v2

@@ 9 Ablation Study: ResearchTown is Robust
We conduct ablation studies on both hyperparameters and model selection. The results show that ResearchTown consistently produces high-quality simulations across a range of settings, demonstrating strong robustness. Detailed experimental configurations are provided in Appendix §I.
Figure 6: Fine-grained similarity evaluation using LLM-as-a-judge for paper writing simulation. We use GPT-4o as the evaluator, prompting it to score each dimension on a scale from 0 to 10. The first five dimensions assess specific aspects of similarity, while the final score (overall similarity) represents an overall score as judged by the LLM.
@@
Ablation on paper number. In paper writing tasks, users can freely assign papers to simulate non-existent work, making robustness to the number of papers essential. As shown in Figure 5, papers cited in the related work section have the greatest positive impact, increasing the similarity score from 66.4 to 66.7 compared to using all papers. In contrast, using only papers cited in the introduction lowers the score to 65.2, while including papers from other sections reduces it further to 58.4. These results highlight the importance of selecting informative references when generating papers. In review writing, the number of papers is fixed, so no ablation study on the paper number is applicable.
Ablation on agent number. For ResearchTown simulation, users can assign different numbers of agents, making robustness to agent number critical for ResearchTown. In Figure 5, in the paper writing task, increasing the agent number improves simulation quality under the agent-aggregation setting. The most notable gain occurs when increasing from 1 to 2, boosting the similarity score from 49.0 to 52.7. Similar trends hold in review writing (Figure 5), where increasing the agent number consistently enhances output quality. The strength score improves from 50.8 to 51.5 when increasing the reviewer from 1 to 5.
Ablation on generation models. The choice of LLMs significantly impacts simulation quality. In addition to GPT-4o-mini, we evaluate two models from different families: Qwen-2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and Deepseek-v3666We point DeepSeek-V3-0324 for use.. In Table 3, for both paper writing tasks, global aggregation (ResearchTown) consistently yields the highest similarity scores across all models. It also achieves the best review difference scores for Qwen-2.5-7B-Instruct and Deepseek-v3. The only exception is GPT-4o-mini, which shows an unexpected increase in review difference under AGG-global. Overall, Deepseek-v3 outperforms GPT-4o-mini, which in turn outperforms Qwen-2.5-7B-Instruct—consistent with their relative performance on other tasks.
Ablation on embedding models. Similarity scores can be computed using different models, and voyage-3777https://blog.voyageai.com/2024/09/18/voyage-3/ serves as an alternative to the text-embedding-3-large used in our main experiments. As shown in Figures 5, 5, and 5, voyage-3 produces consistent trends in ablation studies involving the number of papers and agents. This consistency suggests that ResearchTown is robust to the choice of embedding model, and different models lead to the same conclusions.
10 Discussion: ResearchTown is Effective
Besides computing embedding-based similarities, we provide more types of evaluations here. First, we prompt LLMs to calculate fine-grained similarity scores that assess consistency between real-world data and simulated ones across various dimensions. Next, we evaluate the intrinsic quality of the simulated outputs themselves and compare them with real-world data. Finally, we report results from human evaluations to validate the alignment between LLM-based evaluation and human judgments. More details about LLM-based evaluation are available in Appendix §G, and details about human evaluation are available in Appendix §H.
Automatic evaluation on fine-grained similarity. A high cosine similarity score alone can mask important issues in simulated results. To capture a more complete picture of similarity, we move beyond a single score and instead evaluate across five fine-grained dimensions: topic consistency, method consistency, factual consistency, claim consistency, and application context consistency. These dimensions collectively reflect subcomponents of overall semantic similarity. For evaluation, we use GPT-4o to assign scores from 0 to 10 for each dimension for each paper. As shown in Figure 6, our proposed global aggregation method (ResearchTown) consistently outperforms all other aggregation baselines across these dimensions. This demonstrates that ResearchTown provides a more effective simulation of research activities compared to baselines.
Table 3: Comparison of simulation results with different generation models. For Qwen, we refer to Qwen-2.5-7B-Instruct. For GPT, we refer to GPT-4o-mini. For DS, we refer to Deepseek-v3. For paper writing metrics, we utilize the overall similarity. For review writing metrics, we use $\Delta\mathbf{S}$ to represent its review alignment with the real world.
AGG Type Paper Writing Review Writing
Qwen GPT DS Qwen GPT DS
AGG-self 46.45 46.08 48.62 1.36 1.27 1.11
AGG-agent 53.91 55.24 56.19 1.41 1.19 1.05
AGG-data 65.03 65.30 65.05 1.28 1.26 1.07
AGG-global 65.30 67.51 65.33 0.79 1.51 0.81
Table 4: Evaluation results on novelty and feasibility. Each paper is assigned scores from 0 to 10 for novelty and feasibility. Both LLM-based evaluation and human evaluations are conducted to evaluate the quality of simulated papers. LLM-based evaluation includes results on 1,000 papers, and human evaluation includes results on 40 of them. Simulation represents the outputs of ResearchTown, real-world represents the existing papers.
Evaluation Simulation Real-world
Novelty Feasibility Novelty Feasibility
LLM-based 7.39 6.82 7.85 7.13
Human-based 5.50 7.98 5.90 7.85
@@ 10 Discussion: ResearchTown is Effective
Figure 7: Examples of generated interdisciplinary research papers from ResearchTown. For each example, we include ResearchTown’s responses to two questions: “What is the problem?” and “What are the key components of my approach and results?” as these are the most critical among the five questions mentioned in Appendix §F. Appendix §K provides the full contents of the above two and more examples for interdisciplinary research.
Automatic evaluation on intrinsic quality. In addition to evaluating semantic similarity between simulated and real-world data, we also assess the intrinsic quality of the generated content. Specifically, we focus on two key dimensions: novelty and feasibility, which we consider the two most critical aspects of a research proposal. As shown in Table 4, the simulated outputs still do not match the novelty and feasibility levels of real-world articles but are close to those. This gap indicates that ResearchTown would benefit from a more coordinated agentic workflow to enhance the quality of the generated research outputs.
Human evaluation. Evaluation based on LLMs may introduce bias into the results. To validate the reliability of LLM-based evaluations, we conduct additional human evaluations. For similarity-based assessments, human judgments correlate well with LLM scores, achieving a Pearson correlation of 0.61, indicating reasonable agreement. However, for intrinsic quality evaluations, the correlation between human and LLM scores is low. This is likely due to the inherent ambiguity of such tasks and the need for domain-specific expertise. Despite this, both human and LLM evaluations consistently indicate that simulated papers are slightly less novel than real-world ones—though the gap is relatively small (5.50 vs 5.90 for humans and 7.39 vs 7.85 for LLMs).
11 Case Study: Out-of-distribution Use
As discussed in Section §8, the node masking evaluation in ResearchTown targets in-distribution settings with predefined neighborhoods. In real-world use, however, ResearchTown must generate non-existing papers and reviews without such neighborhoods, requiring automatic construction via paper–researcher matching. This leads to out-of-distribution cases, such as interdisciplinary research, where unrelated papers and researchers form unconventional neighborhoods without prior related works.
ResearchTown can inspire interdisciplinary research. Interdisciplinary research is often challenging due to limited collaboration across fields. ResearchTown addresses this by enabling agents with diverse expertise to read, interact, and co-create novel ideas. For example, as shown in Figure 7, combining NLP and astronomy papers leads to using kinematic models to analyze language evolution, while linking NLP and criminology inspires the use of LLMs to support communities affected by mass incarceration. These domain pairings are rarely explored in existing literature, demonstrating ResearchTown ’s ability to generate innovative, cross-disciplinary research directions.
ResearchTown-written contents might have limited use in the real world. ResearchTown exhibits failure modes when combining too many disparate domains, often producing incoherent or superficial outputs. For example, combining researchers and papers from LLM, biology, criminology, and astronomy, ResearchTown generates a research question of “How does coded language in political discourse influence societal biases, and how can a Bayesian hierarchical model be employed to analyze this effect while simultaneously addressing observational biases in white dwarf population studies?” It simply strings together terminology from different domains without presenting a clear research direction. Such vagueness might hinder the real use of the papers simulated from ResearchTown.
@@ 12 Conclusion
We introduce ResearchTown, a graph-based multi-agent framework that simulates research communities by modeling them as heterogeneous graphs. ResearchTown integrates key research activities—paper reading, writing, and reviewing—into a unified TextGNN-driven inference process. It enables realistic and robust simulations through agent collaboration and facilitates rare interdisciplinary interactions. ResearchTown offers a valuable platform for studying research dynamics and developing algorithms to support automated scientific discovery.
Impact Statement
ResearchTown presents an LLM-based simulation framework that models human research communities as graph-based multi-agent systems, enabling the study of collaboration, knowledge diffusion, and institutional dynamics. By formalizing how agents create, refine, and evaluate academic papers, the simulator can inform the design of autonomous research systems that assist, rather than replace, human researchers. Potential applications include optimizing collaboration structures, identifying systemic bottlenecks in peer review or discovery, and stress-testing scientific workflows under various incentive and communication settings. While the framework is primarily a research tool, we acknowledge that future extensions involving autonomous agents could raise ethical considerations around authorship, influence, and epistemic trust. Our work highlights the imperative for adaptive ethical frameworks that keep pace with technological capabilities while protecting scholarly values.
Acknowledgments
We sincerely appreciate the support from Amazon grant funding project #120359, ”GRAG: Enhance RAG Applications with Graph-structured Knowledge”, and Meta gift funding project ”PERM: Toward Parameter Efficient Foundation Models for Recommenders”.
@@ Appendix A Individual Contribution
Kunlun Zhu Co-lead, benchmark collection, code writing, paper writing, experiment
@@ Appendix B Ethical Concern
The development and deployment of ResearchTown raises several important ethical considerations that we have carefully addressed in our work. We first discuss how ResearchTown prevents dangerous use, including facilitating plagiarism, producing misleading or low-quality claims, and role-playing human researchers. Furthermore, we discuss the attribution and authorship issues for generated content and discuss the model and data license in our work.
@@ B.1 Potential to facilitate plagiarism
Generative AI’s capabilities for image and text generation can potentially lead to plagiarism in research (Elali & Rachid, 2023). To address this, we have implemented safeguards to ensure responsible usage. ResearchTown is designed as an assistive tool to help researchers gather inspiration for papers and review writing, rather than generating complete, ready-to-use content. By design, ResearchTown ensures that its outputs serve as a starting point for further intellectual effort, rather than a replacement for human researchers.
@@ B.2 Potential to produce misleading or low-quality claims
The motivation of our paper is to simulate research activities and generate preliminary research progress (e.g., papers and reviews that are in their condensed bullet-point summarized format) that can be scrutinized and validated by human researchers, ultimately contributing to the acceleration of the research process. We acknowledge that AI-generated ideas may vary in quality, and therefore, these outputs are not intended for direct dissemination. Instead, they serve as initial, unofficial suggestions that require further experimental validation by human researchers. This approach ensures that only rigorously tested and verified research is presented as final, high-quality work.
@@ B.4 Attribution and authorship
Appendix C Artifact
We list all licenses for the data and models used in our paper in this section.
C.1 Data license
All papers in ResearchBench come from top-tier machine learning conferences (ICLR 2024 and NeurIPS 2024). These papers are publicly available and under the license of CC-BY 4.0, allowing for redistribution and sharing. For the evaluation results of ResearchBench, all inputs and outputs are logged and open for access. Additionally, we keep an accessible record of all supplementary papers referenced during ResearchTown’s inference process. All outputs from ResearchTown are released under the licenses of the papers used for generation.
C.2 Model license
Our work relies on multiple foundation models, including GPT-4o-mini, Qwen-2.5-7B-Instruct, Deepseek-v3, text-embedding-3-large, and voyage-3. Specifically, we use gpt-4o-mini-2024-07-18 accessed via the OpenAI API. We use Qwen-2.5-7B-Instruct-Turbo and Deepseek-v3-0324 via the together.ai  [^8] https://www.together.ai/inference inference API. We utilize the official inference API provided by OpenAI and VoyageAI to use text-embedding-3-large and voyage-3 separately.
The GPT-4o-mini, text-embedding-3-large, and voyage-3 models are closed-source and operate under proprietary licenses. We use them only for academic, and non-commercial purposes and ensure all inputs come from publicly available data, complying with their usage restrictions. By contrast, Qwen-2.5-7B-Instruct is released under the permissive Apache 2.0 license, and Deepseek-v3-0324 is available under the MIT License, allowing for broad academic and research use. We make no modifications to these models and use them as-is via their public APIs.
Appendix D ResearchTown Details
In this section, we provide more explanation and implementation details for ResearchTown simulation algorithm. To achieve better performance and efficiency, we design different prompts for each agent function and make the aggregation process run in parallel.
D.1 ResearchTown aggregation setting implementation
@@
AGG-self. We do not rely on any neighborhood information.
$\triangleright$ Paper writing: The LLM agent without profiles brainstorms independently without referencing any external data or other agents’ ideas. This setting extends Equation 7 to
@@
$\triangleright$ Review writing: The LLM agent without profiles writes the review based solely on the paper itself, without considering additional references or other agents. This setting extends Equation 8 to
@@
AGG-agent. We rely only on agent nodes and exclude data nodes.
$\triangleright$ Paper writing: Multiple LLM agents collaborate by sharing their content and insights to produce the final paper’s content. This setting extends Equation 7 to
@@
$\triangleright$ Review writing: Multiple LLM agents collectively review the paper, sharing their input and critiques to form the final review. This setting extends Equation 8 to
@@
AGG-data. We rely only on data nodes and exclude agent nodes.
$\triangleright$ Paper writing: A single LLM agent without profiles reads and synthesizes information from related data sources to write a paper. This setting extends Equation 7 to
@@
$\triangleright$ Review writing: A single LLM agent without profiles produces a review by reading both the paper and its related data sources, integrating the information to form a comprehensive critique. This setting extends Equation 8 to
@@
AGG-global. We include all neighborhood nodes (both agent and data) during aggregation.
$\triangleright$ Paper writing: Multiple LLM agents produce content in parallel while referencing various data sources. Their aggregated outputs, which incorporate insights from both other agents and data, form the final paper. This setting extends Equation 7 to
@@
$\triangleright$ Review writing: Multiple LLM agents each consider the paper and its related works to write their review. The final review is a combination of these integrated perspectives. This setting extends Equation 8 to
@@
D.2 ResearchTown simulation algorithm implementation
@@
D.3 ResearchTown agent function implementation
@@
The aggregation function for classical GNN in Eq 1 and Eq 2, which is often a pooling or mean operation, is used to condense all neighborhood information into one embedding with the same size as the input. Similarly, our TextGNN layers in Eq 4 and Eq 5, act as an aggregation function similar to classical GNN, producing outputs with controlled textual formats and similar lengths with updated information in the neighborhood nodes by summarizing with LLMs. Therefore, the output length of multiple layers of TextGNN would not increase but would remain approximately the same. We achieve such length control in TextGNN via format control in prompting. We specifically designed prompts to ensure each output adheres to pre-defined constraints. These prompt-controlled constraints ensure stable output lengths at every TextGNN layer, avoiding text length inflation with increasing depth. Each aggregation step condenses and prioritizes critical information, effectively filtering less relevant details.
D.4 ResearchTown future application
Any research-related content—e.g., images, codebases, models, or social media posts—can be represented as nodes in the agent-data graph, with edge types like “cite the paper,” “release model,” or “comment on X post” (examples in Figure 1) defining interactions. By specifying appropriate edge types and agent functions, the framework can be extended to simulate tasks such as code writing, model release, panel discussions, or lectures. While we focus on paper and review writing due to their importance, available real-world data, and simplicity, the framework supports broader applications.
Additionally, ResearchTown can be extended to model social dynamics such as peer pressure, collaborations, and institutional roles via agent-agent relationship edges. Our current implementation already includes role-based dynamics (e.g., leader vs. participant), and we plan to support richer simulations of institutional and reputational factors in future work.
Appendix E ResearchBench Details
In this section, we provide the technical details included in the construction process of ResearchBench. We describe the methodologies used for data collection across its three main components, and we name them as: (1) PaperBench, (2) HighImpactPaperBench, and (3) ReviewBench. Statistically, PaperBench and HighImpactPaperBench focus on a paper writing simulation, which contains 1,000 and 100 tasks, respectively. ReviewBench focuses on review writing simulation and includes 200 tasks.
E.1 Data collection details